OpenStudy (anonymous):

Solve the equation for s. 10=1/3s+1

OpenStudy (anonymous):

how is the equation drawn is it \[10=\frac{ 1 }{ 3s }+1\] or is it\[10=\frac{ 1 }{ 3 }s + 1\]

OpenStudy (anonymous):

10=1/3 s+1

OpenStudy (anonymous):

first one or second one?

OpenStudy (anonymous):

2nd

OpenStudy (anonymous):

do you have any multiple choice answers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

u have to slove for s

OpenStudy (anonymous):

s=27

OpenStudy (anonymous):

o

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i got 26 xD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

if you want to check the answer substitute s for whatever you got and see if it works

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

\[10=\frac{ 1 }{ 3 }27+1\]

OpenStudy (anonymous):

you divide first whats 1/3 of 27

OpenStudy (anonymous):

9

OpenStudy (anonymous):

then you add 1 to 9

OpenStudy (anonymous):

and it works!

OpenStudy (anonymous):

what wouldFind the solution set a2 – 64 = 0 Separate the two values with a comma

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

\[10=\frac{ 1 }{ 3 }27+1\] \[10=9+1\] \[10=10\]